OpenStudy (anonymous):

Find sin(arccsc(20)), where arccsc(theta) = the inverse of csc(theta)

hartnn (hartnn):

do you know this property? \(arccsc(x)=arcsin(1/x)\) use this.

hartnn (hartnn):

@KatClaire Hi, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\) did u get what i have written above ? if not, you can just ask.

OpenStudy (anonymous):

Haha sorry I'm trying to piece it all together. My prof didn't go over this so I'm not quite undrestanding the concept of what arc is?

hartnn (hartnn):

arc here means inverse if sin x = y x= arcsin y read as 'x is sin inverse y'

hartnn (hartnn):

now since sin x = 1/ csc x you have property of inverse trignometric function as : arccsc x = arcsin (1/x) so what will be arccsc(20) = ?? in terms of arcsin ?

OpenStudy (anonymous):

Is it 0.05?

OpenStudy (anonymous):

[-\left( 1\div \left| x \right|\sqrt{x ^{2}-1} \right)\]\[-\left( 1\div \left| x \right|\sqrt{x ^{2}-1} \right)\]

OpenStudy (anonymous):

that being my question... do I just sub in 20? Or am I a moron

hartnn (hartnn):

i thought the question is sin(arccsc(20)) = sin(arcsin(1/20)) =1/20 =0.05 is correct

hartnn (hartnn):

why u wanna substitute 20 there ?

OpenStudy (anonymous):

oooooh okay I get it now. I was just getting messed up with derivatives. Thanks for your help!